A body has a negative charge of 2 micro coloumb if body has 2.5×10^12 protons , then number of electrons present in the body will be

Grade:12

2 Answers

Khimraj
5 years ago
C
Khimraj
5 years ago
Charge of a body is given by 
Q = (p-e)*1.6*10-19
Where p = no of proton and e = no of electron
-2*10-6 = (2.5*1012 – e)*1.6*10-19
So e = 1.5*1013
Hope it clears. If you like answer then please approve it.